OpenStudy (anonymous):

what are the real or imaginary solutions of each polynomial equation? x^4-20x^2=-64

OpenStudy (anonymous):

x^4-20x^2+64=0 We can use quadratic techniques to solve this. (x^2)^2-20(x^2)+64 x^2 = (20 +/- sqrt(400-256))/2 x^2 = (20 +/- sqrt(144))/2 = (20 +/- 12)/2 = 10 +/- 6 = 4 or 16 x^2=-2 or x^2=16 x=+/- sqrt(4) or +/- sqrt(16) x=+/- 2 or +/- 4 Factored, the equation is (x^2+2)(x^2-16)=0. You can plug in the values of x to check for extraneous solutions, which you would exclude from the solution list.